For the previous question, if it actually takes 25 minutes to reach the cruising altitude and speed, how much energy is lost to air resistance? Give an answer in GJ (gigajoules, 1 GJ = 10° J) to two significant figures of accuracy.


1
Expert's answer
2021-06-10T10:22:14-0400

From Q1 and Q2 we obtained the values for energy and time in ideal conditions respectively.

Actually, however, it takes 25 minutes. So, the real energy is


"E_R=P_\text{engines}t_R=4·(56·10^6)(25·60)=336·10^9\text{ J}."

Or 340 GJ. So, subtract the work in ideal conditions (Q1) from this value:


"E_\text{lost}=E_R-E_I=340-E_I."